a) Calculate the potential energy of a charge +6nC placed at 0.07m from a charge +76 x10-6C. If the same charge if at 0.07m from a charge -88 x10-6C, find its potential energy.

b) Find the distance from a charge -8 x10-6C to a charge -4 x10-9C if there is a potential energy of 9.5kJ between them. Find the initial force on charge -4 x10-9C.



1
Expert's answer
2021-10-18T11:13:37-0400

a) The potential energy of two point charges is given as follows:


"W = k\dfrac{q_1q_2}{r}"

where "q_1,q_2" are the charges, "r" is the distance between them, "k = 9\times 10^9 Nm^2/C^2" is the Coulomb constant. Substituting the numbers from the task, obtain:

1.


"W =9\times 10^9\cdot \dfrac{6\times10^{-9}\cdot 76\times10^{-6}}{0.07} \approx 0.06J"

2.


"W =9\times 10^9\cdot \dfrac{6\times10^{-9}\cdot (-88)\times10^{-6}}{0.07} \approx -0.07J"


b) Expressing the distance from the first formula, find:


"r = \dfrac{kq_1q_2}{W}"

Substituting the numbers, get:


"r = \dfrac{9\times 10^9\cdot (-8)\times 10^{-6}\cdot (-4)\times 10^{-9}}{9.8\times 10^3} \approx 2.9\times 10^{-8}m"

The force is given by the Coulomb's law:


"F = k\dfrac{q_1q_2}{r^2} = \dfrac{W}{r} = \dfrac{W^2}{kq_1q_2}\\
F = \dfrac{(9.8\times 10^3)^2}{9\times 10^9\cdot (-8)\times 10^{-6}\cdot (-4)\times 10^{-9}} \approx 3.3\times 10^{11}N"

Answer. a) 0.06J and -0.07J. b) "2.9\times 10^{-8}m" and "3.3\times 10^{11}N".